OpenStudy (anonymous):

A sound wave is modeled with the equation y= 1/4 cos 2pi/3 theta. A. Find the period. Explain your method. B. Find the amplitude. Explain your method. C. What is the equation of the midline? What does is represent?

OpenStudy (anonymous):

\[y=\frac{ 1 }{ 4} \cos \frac{ 2\pi }{ 3 } \]theta

OpenStudy (kropot72):

In the equation \[\large y=A \cos(Bx)\] the period T is given by \[\large T=\frac{2\pi}{B}\] and the amplitude is absolute value of A = |A| You need to plug the values into the above to find the values of the period and the amplitude.

OpenStudy (kropot72):

In your equation \[\large B=\frac{2\pi}{3}\] and |A| = 1/4

OpenStudy (anonymous):

So the period is T= 2pi/3 and the amplitude is 1/4?

OpenStudy (kropot72):

The amplitude is 1/4. The period is given by \[\large T=\frac{2\pi}{\frac{2\pi}{3}}\]

OpenStudy (anonymous):

So that's the period, the amplitude is 1/4. How do we find the midline cuz i have to show my work

OpenStudy (kropot72):

\[\large T=\frac{2\pi}{\frac{2\pi}{3}}=\frac{2\pi}{1}\times\frac{3}{2\pi}=?\]

OpenStudy (anonymous):

3?

OpenStudy (kropot72):

Correct, the period is 3 units. There is no y-shift, therefore the equation of the midline is y = 0

OpenStudy (kropot72):

The equation y = 0 is simply the x-axis.

OpenStudy (anonymous):

How do we find the midline though? @kropot72

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (anonymous):

@Astrophysics

OpenStudy (kropot72):

As I posted previously "There is no y-shift, therefore the equation of the midline is y = 0" "The equation y = 0 is simply the x-axis."

OpenStudy (kropot72):